OpenStudy (anonymous):

Help please on properties of exponents? will give a medal! A rectangle has a length of the cube root of 81 inches and a width of 3 to the 2 over 3 power inches. Find the area of the rectangle. a. 3 to the 2 over 3 power inches squared b. 3 to the 8 over 3 power inches squared c. 9 inches squared d.9 to the 2 over 3 power inches squared 2.)Explain how the Quotient of Powers was used to simplify this expression. 2 to the fifth power, over 8 = 2 to the 2nd power a.By finding the quotient of the bases to be one fourth and cancelling common factors b. By finding the quotient of the bases to be one fourth and simplifying the expression c. By simplifying 8 to 23 to make both powers base two and subtracting the exponents d. By simplifying 8 to 23 to make both powers base two and adding the exponents 3.)the cube root of 2 to the seventh power a. 2 to the 3 over 7 power b. 2 to the 7 over 3 power c. 2^21 d. 2^4

OpenStudy (mertsj):

For the first one use the fact that 81 = 3^4

OpenStudy (anonymous):

is it b?

OpenStudy (mertsj):

no

OpenStudy (anonymous):

no its c!

OpenStudy (anonymous):

because isnt the square root 0f 81 9?

OpenStudy (mertsj):

\[\sqrt[3]{81}=\sqrt[3]{3^4}=(3^4)^{\frac{1}{3}}=3^{\frac{4}{3}}\]

OpenStudy (anonymous):

oh so it's a?

OpenStudy (mertsj):

\[3^{\frac{4}{3}}\times 3^{\frac{2}{3}}=3^{\frac{4}{3}+\frac{2}{3}}=3^{\frac{6}{3}}=3^2=9\]

OpenStudy (anonymous):

wait, so then i was correct. the answer is C?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thankyou! can you help me with the last two? :)

OpenStudy (mertsj):

Replace 8 with 2^3 and then subtract the exponents.

OpenStudy (mertsj):

For number 3 remember that in a fractional exponent, the top number is the power and the bottom number is the root.

OpenStudy (anonymous):

is number 2 c?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

now i need help with 3 !

OpenStudy (mertsj):

Mertsj Medals 0 For number 3 remember that in a fractional exponent, the top number is the power and the bottom number is the root.

OpenStudy (anonymous):

a?

OpenStudy (mertsj):

What is the power given in the problem?

OpenStudy (anonymous):

7

OpenStudy (mertsj):

in a fractional exponent, the top number is the power and the bottom number is the root.

OpenStudy (mertsj):

So you chose a which says the exponent is 3 over 7. In the fraction 3/7 is the top number 7?

OpenStudy (anonymous):

oh! the power is 3

OpenStudy (mertsj):

If the exponent is 3/7, then the power is 3. Does your problem say the power is 3?

OpenStudy (anonymous):

no it doesn't the power is 2 in my problem, correct?

OpenStudy (anonymous):

Oh wait, is the answer b?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Thankyou :)

OpenStudy (mertsj):

yw